# Do We Really Need Permutations? Impact of Width Expansion on Linear Mode Connectivity

### v1

@@ Abstract
Recently, Ainsworth et al. (2023) empirically demonstrated that, given two independently trained models, applying a parameter permutation that preserves the input–output behavior allows the two models to be connected by a low-loss linear path. When such a path exists, the models are said to achieve linear mode connectivity (LMC). Prior studies, including Ainsworth et al. (2023), have reported that achieving LMC requires not only an appropriate permutation search but also sufficiently wide models (e.g., a 32 $\times$ width multiplier for ResNet-20). This is broadly believed to be because increasing the model width ensures a large enough space of candidate permutations, increasing the chance of finding one that yields LMC. In this work, we empirically demonstrate that, even without any permutations, simply widening the models is sufficient for achieving LMC when using a suitable softmax temperature calibration. We further explain why this phenomenon arises by analyzing intermediate layer outputs. Specifically, we introduce layerwise exponentially weighted connectivity (LEWC), which states that the output of each layer of the merged model can be represented as an exponentially weighted sum of the outputs of the corresponding layers of the original models. Consequently the merged model’s output matches that of an ensemble of the original models, which facilitates LMC. To the best of our knowledge, this work is the first to show that widening the model not only facilitates nonlinear mode connectivity, as suggested in prior research, but also significantly increases the possibility of achieving linear mode connectivity.
@@ 1 Introduction
Large neural networks (NNs) are widely used across various domains (Vaswani et al., 2017; van den Oord et al., 2016; Zhao et al., 2023), and optimizing their parameters constitutes a massive non-convex optimization problem. Remarkably, stochastic gradient descent (SGD), which is widely employed for NN training, is known to find good solutions despite its simplicity. One hypothesis proposed to explain this seemingly counterintuitive phenomenon is that the landscape of the loss function may be far simpler than previously thought. Several studies (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have reported that different NN solutions can be connected through simple nonlinear paths with almost no increase in loss. Recently, Entezari et al. (2022) conjectured that, after accounting for all permutation symmetries in neural networks, 1.1 may hold.
@@
Here, the barrier refers to the amount of loss increase observed when linearly interpolating the weights between two models. When the barrier between two models is sufficiently small, they are said to exhibit linear mode connectivity (LMC) (Frankle et al., 2020). 1.1 claims that many SGD solutions can be mapped into the same loss basin by applying an appropriate permutation. Indeed, several studies (Ainsworth et al., 2023; Singh & Jaggi, 2020) have experimentally demonstrated the validity of this conjecture across a variety of datasets and models by employing effective permutation-finding techniques such as weight matching, straight-through estimators, and activation matching.
@@
In previous studies (Entezari et al., 2022; Ainsworth et al., 2023; Singh et al., 2024; Ito et al., 2025a; b; Ferbach et al., 2024), it has been widely believed that sufficiently widening the model is required to find a permutation under which the the LMC holds. Intuitively, if the model is not sufficiently wide, the number of candidate permutations becomes limited, making it difficult to discover an appropriate permutation that transfers the models into the same loss basin. For example, Ainsworth et al. (2023) empirically demonstrated that, without increasing ResNet-20’s width by $32\times$ or VGG-16’s by $4\times$, permutations fail to sufficiently reduce the loss barrier on CIFAR-10 dataset. In addition, several studies have shown that widening ResNet-50 trained on the ImageNet dataset improves the test accuracy of the merged model when interpolating the weights of two trained models with permutations. The original conjecture (1.1) also suggests that there may be no permutation that enables LMC unless the model is sufficiently wide.
@@
In this paper, we empirically show that once the model is widened enough, simply averaging the weights of independently trained models without applying any permutation achieves test accuracy comparable to that of the original models. This finding implies that even without aligning models to the same loss basin via permutation, sufficiently widening the models naturally places them within the same basin in terms of test accuracy. Prior works (Nguyen, 2019; Shevchenko & Mondelli, 2020) have demonstrated that increasing model width facilitates the existence of nonlinear low-loss paths between trained models. However, to the best of our knowledge, no study has suggested that it also facilitates connectivity through linear paths. Our results suggest that widening the model itself may play a more critical role in achieving LMC than increasing the number of possible permutations.
@@
Understanding the principle behind LMC is important not only for theoretical interest, such as explaining why SGD performs effectively in deep learning, but also for practical applications like model merging where independently trained models are combined. Several previous studies (Singh & Jaggi, 2020; Wang et al., 2020; Guerrero-Peña et al., 2023) have proposed techniques for model merging, federated learning, and continual learning by leveraging permutation symmetries in neural networks. Weight averaging is already known to work well for models that have been fine-tuned from a shared foundation model. By exploring the principles of LMC without permutations, our study suggests that a similar strategy may also be feasible for models that are trained completely independently, which has been considered difficult to achieve.
@@
We further show that the low-rank structure of weight matrices plays a crucial role in achieving LEWC. Previous work on permutation-based model merging has pointed out that low-rankness of weights is essential for LMC (Ito et al., 2025a; Yunis et al., 2024). We demonstrate that this requirement also applies to LEWC by conducting experiments that vary the degree of weight decay. These results suggest that the possibility for LMC to hold depends strongly on properties of the solutions obtained by SGD.
@@ 2.2 Linear Mode Connectivity (LMC)
Definition 2.1 (Loss Barrier  (Entezari et al., 2022)).
@@ 2.3 Permutation Symmetry and LMC
Ainsworth et al. (2023) showed that two independently trained wide networks can achieve LMC by permutation symmetries through weight matching (WM), which finds permutations that minimize their $L^{2}$ distance. Intuitively, if $\bm{\theta}_{a}\approx\pi(\bm{\theta}_{b})$ for some $\pi$, then $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$ can be seen as almost the same parameters, so their interpolation $\lambda\bm{\theta}_{a}+(1-\lambda)\pi(\bm{\theta}_{b})$ should preserve performance, where $\lambda$ is the merging (interpolation) coefficient. This view suggests that models scattered in parameter space may lie in a common loss basin once permutations are accounted for. However, prior works (Ainsworth et al., 2023; Entezari et al., 2022; Guerrero-Peña et al., 2023; Ito et al., 2025a) indicate that WM requires very large widths (e.g., $32\times$ for ResNet-20, $4\times$ for VGG-16). In contrast, our results show that simply widening the models already improves merged performance monotonically, eventually matching the originals without permutation.
@@ 3 Width Expansion Facilitates High-Accuracy Model Merging
Ainsworth et al. (2023) experimentally demonstrated that in permutation-based model merging, sufficiently wide models are required for LMC to hold. In this section, we empirically show that even without permutations, widening the models improves the performance of merged models. First, we confirm that widening increases test accuracy and reduces test loss. Next, we demonstrate that applying random permutations does not degrade performance. These results suggest that, once the models are sufficiently wide, permutations are no longer essential for achieving LMC.
@@ 3.1 Test Accuracy and Loss
Figure 2: Test losses of merged models without permutations. Figure 2(a) shows the original loss values, while Figure 2(b) shows the values obtained by applying temperature scaling (inverse temperature).
Figure 1 shows the mean and standard deviation of test accuracies over three independent merges of two models. [^1] Because LMC holds only when the models are sufficiently widened, our experiments are limited to relatively simple datasets such as CIFAR-10 and MNIST. In Section E.1, we also provide results on a more complex dataset, CIFAR-100, confirming that widening the model similarly removes the barrier. For comparison, we also include results obtained with the permutations discovered by weight matching (WM) in Figure 12. Details of the models, hyperparameters, and permutation search methods are provided in Appendix B. As shown in the figure, widening the models enables merged models to achieve accuracy comparable to the original ones, even without permutations.
Figure 2 presents the test losses of merged models. As observed in Figure 2(a), the test loss does not decrease sufficiently to match that of the original models, although Figure 1 shows that test accuracy monotonically increases with width. This discrepancy arises because certain transformations of the output distribution can change the loss without affecting accuracy. A concrete example is scaling the logits by an inverse temperature, which alters the cross-entropy loss but leaves the predicted labels unchanged. Since accuracy is the primary objective in classification, it is reasonable to evaluate the loss under an optimal inverse temperature. Accordingly, we estimated the inverse temperature using 20% of the test set and computed calibrated losses on the remaining 80%. The calibrated results are shown in Figure 2(b). With this adjustment, the loss barrier approaches zero as width increases, as expected. In this sense, when we state that LMC holds with a zero loss barrier, we also include the case where the softmax is calibrated with an inverse temperature.
@@ 3.2 Random Permutations
As discussed above, sufficiently wide models achieve performance comparable to the originals even without permutations. This suggests that permutations are not essential once the models are sufficiently wide. To test this further, we applied random permutations before merging two models at $\lambda=1/2$. As shown in Figure 13, the merged models still maintain high accuracy, indicating that permutations are not critical when models are sufficiently wide.
@@ 4.1 Layerwise Exponentially Weighted Connectivity
Zhou et al. (2023) introduced layerwise linear feature connectivity (LLFC), a concept related to LEWC. LLFC explains why LMC holds for permutation-based methods and spawning. LLFC states that, for each layer, the output of the merged model can be expressed as a weighted average of the outputs of the two original models. As sufficient conditions for LLFC, Zhou et al. (2023) proposed weak additivity for ReLU activations and a commutativity property. In our setting, we show in Appendix C that the commutativity property does not hold. This is why, in this work, we introduce LEWC as a concept distinct from LLFC.
@@ 5 Why Does LEWC Emerge in the Wide-Width Regime?
In this section, we clarify why widening the model leads to the satisfaction of LEWC. First, in Section 5.1, we introduce, as sufficient conditions for LEWC, (1) weak additivity for ReLU activations and (2) reciprocal orthogonality. When (1) and (2) are satisfied, LEWC holds. Therefore, in Sections 5.2 and 5.3, we empirically verify whether conditions (1) and (2) are satisfied, and further reveal that the reasons for their validity are due to the low-rank structure of the weights. In other words, conditions (1) and (2) do not hold when the weights are not low-rank, in which case neither LEWC nor LMC holds. Previous studies (Ito et al., 2025a; Galanti et al., 2022) have pointed out that weakening weight decay during SGD training increases the rank of the weights. Therefore, in Section 5.4, we empirically demonstrate that by indirectly increasing the rank of the weight matrices through weakening weight decay, both LEWC and LMC do not hold. This clarifies that the rank of the weight matrices strongly influences the realization of LEWC and LMC.
@@ 5.1 Sufficient Conditions for LEWC
Definition 5.1 (Weak Additivity for ReLU Activations (Zhou et al., 2023)).
@@
The proof of Theorem 5.3 is shown in Section D.1. Regarding the assumption that biases can be ignored, in models such as ResNet and VGG that employ batch normalization after every convolutional layer, the effect of biases can indeed be considered negligible. Moreover, the exponential decay in activation norms is compensated for by the normalization step. Therefore, from Theorem 5.3, LMC can be achieved when Definition 5.1 and Definition 5.2 are satisfied. Thus, in the following two subsections, we empirically confirm whether these two conditions actually hold.
@@ 5.2 Empirical verification of weak additivity for ReLU activations
The Gaussian argument yields a cosine similarity of about 0.93 can be achieved when the dimension is sufficiently large; however, Figure 4 shows cosine similarities higher than 0.93 for wide models, so dimensionality alone does not fully explain the approximate linearity of ReLU. An additional reason is the low-rank structure of weight matrices induced by widening. To illustrate the basic idea, let $\tilde{\bm{z}}_{\ell}^{(a)}$ and $\tilde{\bm{z}}_{\ell}^{(b)}$ be the pre-activations at some layer $\ell$ in the two models. Let $d_{\ell}$ be the dimensionality at this layer, and for simplicity assume $d_{\ell}$ is even. Suppose that in $\tilde{\bm{z}}_{\ell}^{(a)}$, all coordinates except the first $d_{\ell}/2$ are zero (i.e., $\forall i\in\{d_{\ell}/2+1,d_{\ell}/2+2,\dots,d_{\ell}\},\tilde{\bm{z}}_{\ell,i}^{(a)}=0$), and in $\tilde{\bm{z}}_{\ell}^{(b)}$, all coordinates except the last $d_{\ell}/2$ are zero (i.e., $\forall i\in\{1,2,\dots,d_{\ell}/2,\},\tilde{\bm{z}}_{\ell,i}^{(b)}=0$). Then we can show that $\sigma(\tilde{\bm{z}}_{\ell}^{(a)}+\tilde{\bm{z}}_{\ell}^{(b)})=\sigma(\tilde{\bm{z}}_{\ell}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell}^{(b)})$ because, for any $i$, $\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)}+\tilde{\bm{z}}_{\ell,i}^{(b)})=\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)})$ if $i\leq d_{\ell}/2$, and $\sigma(\tilde{\bm{z}}_{\ell,i}^{(b)})$ otherwise. On the other hand, a similar relation also holds for $\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell,i}^{(b)})$. Thus, in this case, ReLU function behaves linearly.
@@
Figure 5: Histogram of standard deviations of the ReLU inputs in the second hidden layer, relative to zero (i.e. $\sqrt{\mathbb{E}\tilde{\bm{z}}_{\ell,i}^{2}}$). Here, we present results for MLPs with width $\times 16$, a VGG-11 scaled $\times 16$, and a ResNet-20 scaled $\times 32$. Most dimensions are concentrated in the leftmost bin, indicating that only a few dimensions are active. Results for all layers are shown in Section E.4.
@@
Figure 6: Proportion of input dimensions with small standard deviation (“Small input std.”) and the proportion of non-overlapping high-variance dimensions (“Non-overlap”) between two trained models for the second ReLU layer. Here, “Small input std.” refers to dimensions whose standard deviation is smaller than one hundredth of the maximum standard deviation across all dimensions, which corresponds to the leftmost bin in Figure 5. As model width increases, the fraction of small-variance dimensions grows and the overlap decreases.
@@
From the above considerations, for the pre-activations at each layer of the two models, having a smaller overlap among the dimensions whose magnitudes deviate significantly from zero makes the linearity of ReLU more likely to hold. In particular, this is more likely when each weight has low rank. This is because one model’s pre-activation is given by $\tilde{\bm{z}}_{\ell}^{(a)}=\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}^{(a)}$, and if many coordinates of $\tilde{\bm{z}}_{\ell}^{(a)}$ are close to zero regardless of the input, then $\bm{W}_{\ell}^{(a)}$ must have low rank. If the rank of $\bm{W}_{\ell}^{(a)}$ is large, then when regarding $\bm{W}_{\ell}^{(a)}$ as a linear mapping, the dimension of its output space is large, so the number of coordinates near zero necessarily decreases. Ito et al. (2025a) showed that as the model width increases, the rank of each layer’s weight matrix becomes relatively small compared to the width, suggesting that widening the model induces low-rank weights and makes the linearity of ReLU more feasible.
@@
We empirically verified whether widening the model reduces, for each layer, the relative number of dimensions with large variance in the pre-activations with respect to zero rather than around their mean. Figure 5 shows the histogram of pre-activation standard deviations at the second hidden layer, demonstrating that most dimensions have negligible variance. This indicates that only a limited number of dimensions effectively determine the model’s output. Figure 6 also shows that as width increases, the overlap between high-variance dimensions of two models decreases, supporting the approximate linearity of ReLU.

### v2

# Do We Really Need Permutations? Impact of Model Width on Linear Mode Connectivity

@@ Abstract
Recently, Ainsworth_ICLR_2023 empirically demonstrated that, given two independently trained models, applying a parameter permutation that preserves the input–output behavior allows the two models to be connected by a low-loss linear path. When such a path exists, the models are said to achieve linear mode connectivity (LMC). Prior studies, including Ainsworth_ICLR_2023, have reported that achieving LMC requires not only an appropriate permutation search but also sufficiently wide models (e.g., a 32 $\times$ width multiplier for ResNet-20). This is broadly believed to be because increasing the model width ensures a large enough space of candidate permutations, increasing the chance of finding one that yields LMC. In this work, we empirically demonstrate that, even without any permutations, simply widening the models is sufficient for achieving LMC when using a suitable softmax temperature calibration. We further explain why this phenomenon arises by analyzing intermediate layer outputs. Specifically, we introduce layerwise exponentially weighted connectivity (LEWC), which states that the output of each layer of the merged model can be represented as an exponentially weighted sum of the outputs of the corresponding layers of the original models. Consequently the merged model’s output matches that of an ensemble of the original models, facilitating LMC. To the best of our knowledge, this work is the first to show that widening the model not only facilitates nonlinear mode connectivity, as suggested in prior research, but also significantly increases the possibility of achieving linear mode connectivity.
@@ 1 Introduction
Large neural networks (NNs) are widely used across various domains (Vaswani_NIPS_2017; Aaron_arxiv_2016; Wayne_arxiv_2023), and optimizing their parameters constitutes a massive non-convex optimization problem. Remarkably, stochastic gradient descent (SGD), which is widely employed for NN training, is known to find good solutions despite its simplicity. One hypothesis proposed to explain this seemingly counterintuitive phenomenon is that the landscape of the loss function may be far simpler than previously thought. Several studies (Garipov_NIPS_2018; Draxler_ICML_2018; Freeman_ICLR_2017) have reported that different NN solutions can be connected through simple nonlinear paths with almost no increase in loss. Recently, Entezari_arxiv_2022 conjectured that, after accounting for all permutation symmetries in NNs, 1.1 may hold.
@@
Here, the barrier refers to the amount of loss increase when linearly interpolating between the weights of the two models. When the barrier between them is sufficiently small, they are said to exhibit linear mode connectivity (LMC) (frankle_ICML_2020). 1.1 claims that many SGD solutions can be mapped into the same loss basin by applying an appropriate permutation. Indeed, several studies (Ainsworth_ICLR_2023; Singh_NIPS_2020) have experimentally validated this conjecture across a variety of datasets and models by employing effective permutation-finding techniques.
@@
In previous studies (Entezari_arxiv_2022; Ainsworth_ICLR_2023; Signh_HiLD_2024; Ito_ICLR_2025; Ito_ICML_2025; Ferbach_AISTATS_2023), it has been widely believed that sufficiently widening the model is required to find a permutation under which LMC holds. Intuitively, if the model is not sufficiently wide, the number of candidate permutations is limited, making it difficult to discover an appropriate permutation that brings the models into the same loss basin. For example, Ainsworth_ICLR_2023 empirically demonstrated that, without increasing ResNet-20’s width by $32\times$ or VGG-16’s by $4\times$, permutations fail to sufficiently reduce the loss barrier on the CIFAR-10 dataset. In addition, several studies have shown that widening ResNet-50 trained on the ImageNet dataset improves the test accuracy of the merged model when interpolating the weights of two trained models with permutations (Ainsworth_ICLR_2023). The original conjecture (1.1) also suggests that there may be no permutation that enables LMC unless the model is sufficiently wide.
@@
In this paper, we empirically show that once the model is sufficiently wide, simply averaging the weights of independently trained models without applying any permutation achieves test accuracy comparable to that of the original models. This finding implies that even without aligning models to the same loss basin via permutation, sufficiently widening the models naturally places them within the same basin in terms of test accuracy. Prior works (Nguyen_ICML_2019; Shevchenko_ICML_2020) have demonstrated that increasing model width facilitates the existence of nonlinear low-loss paths between trained models. However, to the best of our knowledge, no study has suggested that it also facilitates connectivity through linear paths. Our results suggest that widening the model itself may play a more critical role in achieving LMC than increasing the number of possible permutations.
@@
Understanding the principle behind LMC is important not only for theoretical reasons, such as explaining the effectiveness of SGD in deep learning, but also for practical applications like model merging. Prior work (Singh_NIPS_2020; Wang_ICLR_2020; Pena_CVPR_2023) has leveraged permutation symmetries in neural networks for model merging, federated learning, and continual learning. While weight averaging is known to work well for models fine-tuned from a shared foundation model, merging independently trained models remains challenging. By studying LMC without permutations, we suggest that a similar strategy may also be feasible in this setting.
@@
We further show that the low-rank structure of weight matrices plays a crucial role in achieving LEWC. Previous work on permutation-based model merging has pointed out that low-rankness of weights is essential for LMC (Ito_ICLR_2025; Yunis_arxiv_2024). We demonstrate that this requirement also applies to LEWC by conducting experiments that vary the degree of weight decay. These results suggest that the possibility for LMC to hold depends strongly on properties of the solutions obtained by SGD.
@@ 2.2 Linear Mode Connectivity (LMC)
Definition 2.1 (Loss Barrier  (Entezari_arxiv_2022)).
@@ 2.3 Permutation Symmetry and LMC
Ainsworth_ICLR_2023 showed that two independently trained wide networks can achieve LMC by permutation symmetries through weight matching (WM), which finds permutations that minimize their $L^{2}$ distance. Intuitively, if $\bm{\theta}_{a}\approx\pi(\bm{\theta}_{b})$ for some $\pi$, then $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$ can be seen as almost the same parameters, so their interpolation $\lambda\bm{\theta}_{a}+(1-\lambda)\pi(\bm{\theta}_{b})$ should preserve performance, where $\lambda$ is the merging (interpolation) coefficient. This view suggests that models scattered in parameter space may lie in a common loss basin once permutations are accounted for. However, prior works (Ainsworth_ICLR_2023; Entezari_arxiv_2022; Pena_CVPR_2023; Ito_ICLR_2025) indicate that WM requires very large widths (e.g., $32\times$ for ResNet-20, $4\times$ for VGG-16). In contrast, our results show that simply widening the models already improves merged performance monotonically, eventually matching the originals without permutation.
@@ 3 Width Expansion Facilitates High-Accuracy Model Merging
Ainsworth_ICLR_2023 experimentally demonstrated that, in permutation-based model merging, sufficiently wide models are required for LMC to hold. In this section, we empirically show that even without permutations, widening the models improves the performance of merged models. First, we confirm that widening increases test accuracy and reduces test loss. Next, we demonstrate that applying random permutations does not degrade performance. These results suggest that, once the models are sufficiently wide, permutations are no longer essential for achieving LMC.
@@ 3.1 Test Accuracy and Loss
Figure 2: Test losses of merged models without permutations. Figure 2(a) shows the original loss values, while Figure 2(b) shows the values obtained by applying temperature scaling.
Figure 1 shows the mean and standard deviation of test accuracies over three independent merges of two models. [^1] Because LMC holds only when the models are sufficiently widened, our experiments are limited to relatively simple datasets such as CIFAR-10 and MNIST. In Section F.1, we also provide results on a more complex dataset, CIFAR-100, confirming that widening the model similarly removes the barrier. For comparison, we also include results obtained with the permutations discovered by weight matching (WM) in Figure 16. Details of the models, hyperparameters, and permutation search methods are provided in Appendix B. As shown in the figure, widening the models enables merged models to achieve accuracy comparable to the original ones, even without permutations.
Table 1: Barrier values ($\lambda=1/2$) with and without the permutation found by WM. Loss barriers use calibrated losses in both settings. Width multipliers are $16\times$ for MLP, $16\times$ for VGG-11, and $32\times$ for ResNet-20. For sufficiently wide models, barriers are small even without the permutation.
Network Dataset Without perm With perm
Acc. barrier (%) Loss barrier Acc. barrier (%) Loss barrier
MLP MNIST $519\pm 0.225 0.0$$13\pm 0.004 0.0$$27\pm 0.139 0.0$$003\pm 0.003$
FMNIST $2.467\pm 0.685$ $0.056\pm 0.014$ $4.925\pm 4.379$ $0.155\pm 0.111$
VGG-11 CIFAR-10 $1.308\pm 1.594$ $0.066\pm 0.054$ $7.000\pm 3.694$ $0.177\pm 0.039$
ResNet-20 $2.694\pm 1.493$ $0.087\pm 0.051$ $5.135\pm 2.935$ $0.173\pm 0.099$
Figure 2 presents the test losses of merged models. As observed in Figure 2(a), the test loss does not decrease sufficiently to match that of the original models, although Figure 1 shows that test accuracy monotonically increases with width. This discrepancy arises because certain transformations of the output distribution can change the loss without affecting accuracy. A concrete example is scaling the logits by an inverse temperature, which alters the cross-entropy loss but leaves the predicted labels unchanged. Since accuracy is the primary objective in classification, it is reasonable to evaluate the loss under an optimal inverse temperature. Accordingly, we estimated the inverse temperature using 20% of the test set and computed calibrated losses on the remaining 80%. The calibrated results are shown in Figure 2(b). With this adjustment, the loss barrier approaches zero as width increases, as expected. In this sense, when we state that LMC holds with a zero loss barrier, we also include the case where the softmax is calibrated with an inverse temperature. For a more detailed view, Table 1 reports the barrier values separately for the cases with and without applying the permutation. We find that the barriers remain very small even without the permutation.
@@ 3.2 Random Permutations
As discussed above, sufficiently wide models achieve performance comparable to the originals even without permutations. This suggests that permutations are not essential once the models are sufficiently wide. To test this further, we applied random permutations before merging two models at $\lambda=1/2$. As shown in Figure 18(a), the merged models still maintain high accuracy, indicating that permutations are not critical when models are sufficiently wide.
@@ 4.1 Layerwise Exponentially Weighted Connectivity
Zhou_NIPS_2023 introduced layerwise linear feature connectivity (LLFC), a concept related to LEWC. LLFC explains why LMC holds for permutation-based methods and spawning. LLFC states that, for each layer, the output of the merged model can be expressed as a weighted average of the outputs of the two original models. As sufficient conditions for LLFC, Zhou_NIPS_2023 proposed weak additivity for ReLU activations and a commutativity property. In our setting, we show in Appendix C that the commutativity property does not hold. This is why, in this work, we introduce LEWC as a concept distinct from LLFC.
@@ 5 Why Does LEWC Emerge in the Wide-Width Regime?
In this section, we clarify why widening the model leads to the satisfaction of LEWC. First, in Section 5.1, we introduce, as sufficient conditions for LEWC, (1) weak additivity for ReLU activations and (2) reciprocal orthogonality. When (1) and (2) are satisfied, LEWC holds. Therefore, in Sections 5.2 and 5.3, we empirically verify whether conditions (1) and (2) are satisfied, and further reveal that the reasons for their validity are due to the low-rank structure of the weights. In other words, conditions (1) and (2) do not hold when the weights are not low-rank, in which case neither LEWC nor LMC holds. Previous studies (Ito_ICLR_2025; Tomer_arxiv_2023) have pointed out that weakening weight decay during SGD training increases the rank of the weights. Therefore, in Section 5.4, we empirically demonstrate that by indirectly increasing the rank of the weight matrices through weakening weight decay, both LEWC and LMC do not hold. This clarifies that the rank of the weight matrices strongly influences the realization of LEWC and LMC.
@@ 5.1 Sufficient Conditions for LEWC
Definition 5.1 (Weak Additivity for ReLU Activations (Zhou_NIPS_2023)).
@@
For two bias-free models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, if Definitions 5.1 and 5.2 hold, then LEWC is satisfied.
The proof of Theorem 5.3 is shown in Section E.1. Regarding the assumption that biases can be ignored, in models such as ResNet and VGG that employ batch normalization after every convolutional layer, the effect of biases can indeed be considered negligible. Moreover, the exponential decay in activation norms is compensated for by the normalization step. Therefore, from Theorem 5.3, LMC can be achieved when Definition 5.1 and Definition 5.2 are satisfied. Thus, in the following two subsections, we empirically confirm whether these two conditions actually hold.
@@ 5.2 Empirical verification of weak additivity for ReLU activations
The Gaussian argument implies that a cosine similarity of about 0.93 can be achieved when the dimension is sufficiently large; however, Figure 4 shows cosine similarities higher than 0.93 for wide models, so dimensionality alone does not fully explain the approximate linearity of ReLU. An additional reason is the low-rank structure of weight matrices induced by widening. To illustrate the basic idea, let $\tilde{\bm{z}}_{\ell}^{(a)}$ and $\tilde{\bm{z}}_{\ell}^{(b)}$ be the pre-activations at some layer $\ell$ in the two models. Let $d_{\ell}$ be the dimensionality at this layer, and for simplicity assume $d_{\ell}$ is even. Suppose that in $\tilde{\bm{z}}_{\ell}^{(a)}$, all coordinates except the first $d_{\ell}/2$ are zero (i.e., $\forall i\in\{d_{\ell}/2+1,d_{\ell}/2+2,\dots,d_{\ell}\},\tilde{\bm{z}}_{\ell,i}^{(a)}=0$), and in $\tilde{\bm{z}}_{\ell}^{(b)}$, all coordinates except the last $d_{\ell}/2$ are zero (i.e., $\forall i\in\{1,2,\dots,d_{\ell}/2,\},\tilde{\bm{z}}_{\ell,i}^{(b)}=0$). Then we can show that $\sigma(\tilde{\bm{z}}_{\ell}^{(a)}+\tilde{\bm{z}}_{\ell}^{(b)})=\sigma(\tilde{\bm{z}}_{\ell}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell}^{(b)})$ because, for any $i$, $\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)}+\tilde{\bm{z}}_{\ell,i}^{(b)})=\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)})$ if $i\leq d_{\ell}/2$, and $\sigma(\tilde{\bm{z}}_{\ell,i}^{(b)})$ otherwise. On the other hand, a similar relation also holds for $\sigma(\tilde{\bm{z}}_{\ell,i}^{(a)})+\sigma(\tilde{\bm{z}}_{\ell,i}^{(b)})$. Thus, in this case, ReLU function behaves linearly.
@@
Figure 5: Histogram of the square root of the second (uncentered) moment of ReLU inputs in the second hidden layer (i.e., $\sqrt{\mathbb{E}\tilde{\bm{z}}_{\ell,i}^{2}}$). We present results for MLPs with width $\times 16$, a VGG-11 scaled by $\times 16$, and a ResNet-20 scaled by $\times 32$. Most dimensions fall into the leftmost bin, indicating that only a small number of dimensions are active. Results for all layers are provided in Section F.7.
@@
Figure 6: Proportion of input dimensions with a small square root of the second moment (“Small input moment”) and the proportion of non-overlapping large-moment dimensions (“Non-overlap”) between two trained models for the second ReLU layer. Here, “Small input moment” refers to dimensions whose second moment is smaller than one hundredth of the maximum second moment across all dimensions, corresponding to the leftmost bin in Figure 5. As the model width increases, the fraction of small-moment dimensions grows and the overlap decreases.
@@
From the above considerations, for the pre-activations at each layer of the two models, having a smaller overlap among the dimensions whose magnitudes deviate significantly from zero makes the linearity of ReLU more likely to hold. In particular, this is more likely when each weight has low rank. This is because one model’s pre-activation is given by $\tilde{\bm{z}}_{\ell}^{(a)}=\bm{W}_{\ell}^{(a)}\bm{z}_{\ell-1}^{(a)}$, and if many coordinates of $\tilde{\bm{z}}_{\ell}^{(a)}$ are close to zero regardless of the input, then $\bm{W}_{\ell}^{(a)}$ must have low rank. If the rank of $\bm{W}_{\ell}^{(a)}$ is large, then when regarding $\bm{W}_{\ell}^{(a)}$ as a linear mapping, the dimension of its output space is large, so the number of coordinates near zero necessarily decreases. Ito_ICLR_2025 showed that as the model width increases, the rank of each layer’s weight matrix becomes relatively small compared to the width, suggesting that widening the model induces low-rank weights and makes the linearity of ReLU more feasible. [^2] Low rankness does not by itself imply coordinate level sparsity, which is the more direct driver of weak additivity. Nonetheless, our experiments indicate that the two are tightly coupled in practice: varying weight decay changes the rank of the weight matrices, and we observe a corresponding change in coordinate level sparsity of the pre-activations.
@@
We empirically verified whether widening the model reduces, for each layer, the relative number of dimensions with a large second moment of the pre-activations. Figure 5 shows the histogram of the square root of the pre-activation second moment at the second hidden layer, demonstrating that most dimensions have negligible second moment. This indicates that only a limited number of dimensions effectively contribute to the model’s output. Figure 6 further shows that as the width increases, the overlap between the large-moment dimensions of two independently trained models decreases, supporting the approximate linearity of ReLU.